81% of 200 is what number?
Found 2 solutions by jim_thompson5910, rfer:
Answer by jim_thompson5910(35256)


81% of 200 is what number?

81% of 200 is what number

0.81 * 200 = x

0.81 * 200 = x

162 = x

x = 162

---------------------------------------------------

Answer: 81% of 200 is 162

Answer by rfer(16322)

200*0.81=162